OpenStudy (anonymous):

The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Write two equations which represent the two sentences. Solve the system you wrote for this using any technique.

OpenStudy (anonymous):

2x + y = 46 y = x + 4 subtract: 2x = -x + 42 3x = 42 x = 14 y = 14+4 = 18 sides of the triangle are 14,14 and 18